OpenStudy (anonymous):

A mixture of 30 pounds of candy sells for $1.10 a pound. The mixture consists of chocolates worth $1.50 a pound and chocolates worth 90¢ a pound. How many pounds of the $1.50 chocolate were used to make the mixture?

hero (hero):

x + y = 30 1.50x + .90y = 1.10(30)

hero (hero):

^Hint

hero (hero):

@zse, are you working on it or are you stumped?

OpenStudy (anonymous):

im stuck at this question

hero (hero):

I gave you a hint. Have you tried solving the system?

OpenStudy (anonymous):

okay i havent done this in a while can you explain:P

hero (hero):

You haven't done this in a while? You seem like you've never done it. You should at least know the first step.

hero (hero):

Why don't you start by multiplying 1.10(30)

OpenStudy (anonymous):

i did that i got 33

hero (hero):

Oh, I see. I you meant that 1.10(30) = 33

hero (hero):

So now you have x + y = 30 1.50x + .90y = 33

hero (hero):

Now isolate y in both equations

hero (hero):

Let me know what you get.

OpenStudy (anonymous):

ugh i dont know:( i forgot i learned this last year

hero (hero):

If you have x + y = 30, how difficult is it to subtract x from both sides to get y = 30 - x

hero (hero):

Now isolate y for the second one.

hero (hero):

Obviously you have to use algebra to do that. You couldn't possibly have completely forgotten algebra. If you did, you might need to figure out where you hid your brain.

OpenStudy (anonymous):

GASP! my friend well obviously i did your smart and you get this stuff im the type of person who has to see how the process goes:/

OpenStudy (anonymous):

sorry for waisting your time

hero (hero):

To isolate the second equation: 1.50x + .90y = 33, do the following 1. Subtract 1.50x from both sides to get .90y = 33 - 1.50x 2. Divide both sides by .90 to get: \(\large y = \frac{33 - 1.50x}{.90}\) Now you have two equations with y isolated.

hero (hero):

What do you think is the next step?

OpenStudy (anonymous):

subtract 33 - 1.50x?

hero (hero):

How about set y = y ?

hero (hero):

You don't seem to have a clue what is going on.

OpenStudy (anonymous):

no i dont maybe this website isnt for me since i cant figure it out i mean you know what your doing but i just cant figure it out like i said i need to see how the process goes

hero (hero):

Okay, I will show you

hero (hero):

on vyew

OpenStudy (anonymous):

whats vyew

hero (hero):

You'll find out. Click this: http://vyew.com/room#/418625/Open_Study_Math

hero (hero):

Why'd you leave?

OpenStudy (anonymous):

its like a tiny chat?

hero (hero):

It's an online whiteboard

hero (hero):

You said you wanted to see the steps. I will show them to you.

OpenStudy (anonymous):

i got it

OpenStudy (anonymous):

i mean im on the web